a bag contains two black balls, three green balls and four
red balls. A ball is picked from the bag at random. What
is the probability that it is;
1. black
2. green
3. red
4. yellow
5. not black
6. either black or red?

Found 2 solutions by rothauserc, Edwin McCravy:
Answer by rothauserc(4718)

There are nine balls, then
:
1) 2/9
:
2) 3/9 = 1/3
:
3) 4/9
:
4) 0/9 = 0
:
5) 7/9
:
6) 2/9 + 4/9 = 6/9 = 2/3
:

Answer by Edwin McCravy(20055)

a bag contains two black balls, three green balls and four
red balls.
That's a total of 2 blacks + 3 greens + 4 reds = 9 balls

A ball is picked from the bag at random. What
is the probability that it is;
1. black
That's 2 black balls out of 9 balls, so the probability is 2/9.
That doesn't reduce so we leave it 2/9. 

2. green
That's 3 green balls out of 9 balls, so the probability is 3/9.
That reduces to 1/3

3. red
That's 4 red balls out of 9 balls, so the probability is 4/9.
That doesn't reduce so we leave it 4/9. 

4. yellow
There are no yellow balls in the bag.
That's 0 yellow balls out of 9 balls, so the probability is 0/9,
which is just 0.  When something is impossible, its probability
is 0.

5. not black
There are 3+4 or 7 balls that are not black, because 3 are green
and 4 are red.
That's 7 non-black balls out of 9 balls, so the probability is 7/9.
That doesn't reduce so we leave it 7/9.

6. either black or red?
There are 2+4 or 6 balls that are either black or red, because 2 
are black and 4 are red.
That's 6 black or red balls out of 9 balls, so the probability 
is 6/9.  That reduces to 2/3.

Edwin